Formulate the differential equation representing the relationship of current (i), voltage (v), inductance(L) and time t.


1
Expert's answer
2022-02-07T16:19:16-0500

Solution:

The (variable) voltage across the resistor is given by:

"V_{R}=i R"

The (variable) voltage across the inductor is given by:

"V_{L}=L \frac{d i}{d t}"

Kirchhoff's voltage law says that the directed sum of the voltages around a circuit must be zero. This results in the following differential equation:

"R i+L \frac{d i}{d t}=V"

where current (i), voltage (v), inductance(L) and time t.

This is the required D.E.